the length of a rectangle is twice its breadth. If the perimeter is 72 metre, find the length and breadth of the rectangle.
Answer by macston(5194)

W=breadth; L=length=2W P=perimeter=2(L+W)=72m
P=2(L+W)
72m=2(L+W) Divide each side by 2
36m=L+W Substitute for L
36m=2W+W
36m=3W Divide each side by 3.
12m=W ANSWER 1: The breadth is 12 meters.
L=2W=2(12m)=24m ANSWER 2: The length is 24 meters.
CHECK:
P=2(L+W)
72m=2(24m+12m)
72m=2(36m)
72m=72m